Grandpa lunn is twice as old as his son 25 years ago the ratio strange enough was 3:1 when does grandpa lunn turn 100

Answer by ewatrrr(24785)

Grandpa lunn is twice as old as his son 25 years ago the ratio strange enough was 3:1 when does grandpa lunn turn 100

Hi
Let s represent the son's age and G the Granpa's age
Question States***
G = 2s |NOW
&
G-25 = 3(s-25) |25 years ago...
2s - 25 = 3s - 75
50 = s |son's age NOW
G = 2s = 100
Granpa is 100 NOW
CHECKING our answer***
25 years ago...
75 = 3*25 Checks
Wish You the Best in your Studies.